OpenStudy (anonymous):

in the triangle below, calculate the sec 0? A. 5/4 B. 5/3 C. 3/4 D. 4/5

OpenStudy (anonymous):



OpenStudy (anonymous):

@yoyointegrals

OpenStudy (anonymous):

sec = 1/cos

OpenStudy (anonymous):

cos(o) = 4/5

OpenStudy (anonymous):

since we want sec, we invert the equation to be sec(o) = 5/4

OpenStudy (anonymous):

@Luis_Rivera help?

OpenStudy (anonymous):

@katlin95 : Dont ask for answers ,U should learn from examples and then try to figure out the answer by yourself !

OpenStudy (anonymous):

i am learning i just wanna double check my answers..

OpenStudy (anonymous):

@yoyointegrals wouldnt my answer be A?

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

@Eyad u should go else where with your peanut gallery comments